Use a direct proof to show that every odd integer is the difference of two squares.  


1
Expert's answer
2021-12-01T05:05:47-0500

Let "a\in Z" . Let's consider "M=(a+1)^2-a^2=a^2+2a+1-a^2=2a+1" . Formula "2a+1" is the general formula for an odd number, so every odd number can be written as the difference of two squares. Thus, the statement has been proved.